# Lattice point enumeration of polytopes associated to integer compositions

### v1

@@ 1. Introduction
The face lattice of composition polytopes is well understood. As observed in [11, Theorem 20], ${\mathcal{Q}}_{\sigma}$ is combinatorially isomorphic to a product of simplices, the dimensions of which are the parts of $\sigma$. The lattice point enumeration of composition polytopes, which is more interesting, has been studied for the more general classes of arbor polytopes [5] and Pitman–Stanley polytopes [11]. Chapoton defined the $h$-vector of a composition $\sigma$ of $n$ (more generally, of an arbor) as the sequence $h(\sigma)=(h_{i}(\sigma))_{0\leq i\leq n}$, where $h_{i}(\sigma)$ is the number of points in ${\mathcal{Q}}_{\sigma}\cap{\mathbb{N}}^{n}$ having exactly $i$ nonzero coordinates (this vector should not be confused with the $h$-vector of the simple polytope ${\mathcal{Q}}_{\sigma}$). We note that $h_{0}(\sigma)=h_{n}(\sigma)=1$. The following intriguing conjecture has been stated more generally for the class of arbor polytopes [5, Conjecture 0.1]; we refer to the discussion in [5, Section 0.1] for the motivation behind it.
@@
In particular, $h(\sigma)$ should be palindromic, meaning that $h_{i}(\sigma)=h_{n-i}(\sigma)$ for $0\leq i\leq n$. Moreover, it is suggested in [5, Section 0.1] that $h(\sigma)$ should in fact be equal to the $h$-vector of an $n$-dimensional flag simplicial polytope. Thus, in view of Gal’s conjecture [6, Conjecture 2.1.7] (see also [1, Section 3]) on gamma-positivity, it is reasonable to expect that the generating polynomial $h(\sigma,t):=\sum_{i=0}^{n}h_{i}(\sigma)t^{i}$ of $h(\sigma)$ is $\gamma$-positive, meaning that
@@
Our second main result confirms another curious conjecture of Chapoton [5], concerning the Ehrhart polynomial of composition polytopes. We recall (see, for instance, [9, Section 4.6]) that the Ehrhart polynomial ${\rm Ehr}(Q,t)$ of a lattice polytope ${\mathcal{Q}}\subset{\mathbb{R}}_{\geq 0}^{n}$ has the defining property that ${\rm Ehr}({\mathcal{Q}},m)=\#(m{\mathcal{Q}}\cap{\mathbb{N}}^{n})$ for every $m\in{\mathbb{N}}$. The zeta polynomial (see [9, Section 3.12]) ${\mathcal{Z}}(P,t)$ of a finite poset $(P,\preceq)$ has the defining property that ${\mathcal{Z}}(P,m+1)$ is equal to the number of $m$-element multichains in $P$, meaning sequences $(p_{1},p_{2},\dots,p_{m})$ of elements of $P$ such that $p_{1}\preceq p_{2}\preceq\cdots\preceq p_{m}$, for every $m\in{\mathbb{N}}$. Given a composition $\sigma$ of $n$, Chapoton [5, Section 1] has considered the set $P_{\sigma}$ of lattice points in ${\mathcal{Q}}_{\sigma}$ as a subposet of the product poset ${\mathbb{N}}^{n}$.
@@
Theorem 1.2 is proven in Section 2; the proof yields an explicit combinatorial interpretation of the numbers $\gamma_{i}(\sigma)$ which appear in Equation (2). Section 2 also shows that $h(\sigma)=h({\rm rev}(\sigma))$ for every composition $\sigma$. Two proofs of Theorem 1.3, one essentially bijective and another based on the results of [11, Section 3], are given in Section 3 and a combinatorial interpretation of the $h^{\ast}$-polynomial of ${\mathcal{Q}}_{\sigma}$ is deduced via the theory of pure shellability of posets.
Throughout this paper, we set ${\mathbb{N}}=\{0,1,2,\dots\}$ and $n]=\{1,2,\dots,n\}$ for $n\in{\mathbb{N}}$. We denote by #S\#S the cardinality of a finite set $S$ and write $\sigma\models n$ when $\sigma$ is a composition of $n$. Our notation and terminology about posets follows [9, Chapter 3]. Background on Ehrhart theory and face enumeration of simplicial complexes and polytopes can be found in [3, 7] [9, Chapter 4] and [8], respectively.
@@ 3. Ehrhart and Zeta polynomials
Part of our second proof of Theorem 1.3 follows the proof of [11, Equation (33)], which is a formula for the number of lattice points of a Pitman–Stanley polytope. To avoid making too long of a digression, we assume familiarity with Sections 3 and 4 of [11] and only sketch this part of the proof.
@@
Let us consider the poset $\Omega_{\sigma}$ which consists of two disjoint chains $\alpha_{1}\prec\alpha_{2}\prec\cdots\prec\alpha_{n}$ and $\beta_{1}\prec\beta_{2}\prec\cdots\prec\beta_{n}$, together with the additional cover relations $\alpha_{s_{i}}\prec\beta_{s_{i}}$ for every $i\in[k]$. We let $C=\{\beta_{1},\beta_{2},\dots,\beta_{n}\}$ and ${\mathbf{u}}=(u_{1},u_{2},\dots,u_{n})$, where $u_{j}=jm$ for every $j\in[n]$ and where $m\in{\mathbb{N}}$ is fixed. In this situation, Pitman and Stanley [11, Section 3] have defined the order cone section ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ as the polytope of all order preserving maps $f:\Omega_{\sigma}\rightarrow{\mathbb{R}}_{\geq 0}$ such that $f(\beta_{j})=u_{j}$ for every $j\in[n]$. Setting $f_{j}=f(\alpha_{j})$ for every $j\in[n]$ and $f_{0}=0$ we have
@@
for every $i\in[k]$. The transformation $x_{j}=f_{j}-f_{j-1}$ for $j\in[n]$ then shows that ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ is integrally equivalent to $m{\mathcal{Q}}_{\sigma}$. Theorem 5 in [11] gives a formula for the number of lattice points of ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$. By an argument similar to that on [11, p. 617], this formula yields in our situation that
@@
Given a word $w=(w_{1},w_{2},\dots,w_{n})\in[n]^{n}$, we denote by ${\rm asc}(w)$ be the number of indices (called ascents) $i\in[n]$ such that $w_{i-1}<w_{i}$, where $w_{0}:=1$.
@@
The proof employs the basics of the theory of EL-shellability of posets [4] [12, Lecture 3], which we briefly recall. Let $(P,\preceq)$ be a finite graded poset which has a minimum element $\hat{0}$ and a maximum element $\hat{1}$ and let ${\mathcal{E}}(P)$ be the set of covering relations of $P$. An edge labeling of $P$ (with integer labels) is any map $\lambda:{\mathcal{E}}(P)\rightarrow{\mathbb{Z}}$. Let $C$ be an unrefinable chain $p_{0}\prec p_{1}\prec\cdots\prec p_{m}$ of elements of $P$, so that $(p_{i-1},p_{i})\in{\mathcal{E}}(P)$ for every $i\in[m]$. We define the label of $C$ with respect to $\lambda$ as the sequence
@@
and the descent set of $C$ as ${\rm Des}(C)=\{i\in[m-1]:\lambda(p_{i-1},p_{i})>\lambda(p_{i},p_{i+1})\}$. We say that $C$ is rising with respect to $\lambda$ if $\lambda(C)$ is weakly increasing and that $C$ is lexicographically smaller (with respect to $\lambda$) than an unrefinable chain $C^{\prime}$ in $P$ of the same length if $\lambda(C)$ precedes $\lambda(C^{\prime})$ in the lexicographic order. An edge labeling $\lambda$ of $P$ is called an EL-labeling if for every nonsingleton interval $[u,v]$ in $P$: (i) there is a unique rising maximal chain in $[u,v]$; and (ii) this chain is lexicographically smallest among all maximal chains in $[u,v]$, with respect to $\lambda$. Then
@@
where ${\mathcal{M}}(P)$ is the set of all maximal chains of $P$ and ${\rm des}(C)=\#{\rm Des}(C)$ is the number of descents of $C$ for every $C\in{\mathcal{M}}(P)$. Here $\Delta(P)$ stands for the order complex of $P$ (meaning, the simplicial complex consisting of all chains in $P$) and the $h$-polynomial $h(\Delta(P),t)$ is defined as $h(\Delta(P),t)=\sum_{F\in\Delta(P)}t^{\#F}(1-t)^{d+1-(\#F)}$, where $d$ is the rank of $P$.
@@
Remark $3.$
A generalization of Proposition 3.2, valid for the polytopes in ${\mathbb{R}}^{n}$ defined by the equations

### v2

@@ 1. Introduction
The face lattice of composition polytopes is well understood. As observed in [12, Theorem 20], ${\mathcal{Q}}_{\sigma}$ is combinatorially isomorphic to a product of simplices, the dimensions of which are the parts of $\sigma$. The lattice point enumeration of composition polytopes, which is more interesting, has been studied for the more general classes of arbor polytopes [5] and Pitman–Stanley polytopes [12]. Chapoton defined the $h$-vector of a composition $\sigma$ of $n$ (more generally, of an arbor) as the sequence $h(\sigma)=(h_{i}(\sigma))_{0\leq i\leq n}$, where $h_{i}(\sigma)$ is the number of points in ${\mathcal{Q}}_{\sigma}\cap{\mathbb{N}}^{n}$ having exactly $i$ nonzero coordinates (this vector should not be confused with the $h$-vector of the simple polytope ${\mathcal{Q}}_{\sigma}$). We note that $h_{0}(\sigma)=h_{n}(\sigma)=1$. The following intriguing conjecture has been stated more generally for the class of arbor polytopes [5, Conjecture 0.1]; we refer to the discussion in [5, Section 0.1] for the motivation behind it.
@@
In particular, $h(\sigma)$ should be palindromic and unimodal, meaning that $h_{i}(\sigma)=h_{n-i}(\sigma)$ for all $0\leq i\leq n$ and $h_{0}(\sigma)\leq h_{1}(\sigma)\leq\cdots\leq h_{\lfloor n/2\rfloor}(\sigma)$. Moreover, it is suggested in [5, Section 0.1] that $h(\sigma)$ should in fact be equal to the $h$-vector of an $n$-dimensional flag simplicial polytope. Thus, in view of Gal’s conjecture [6, Conjecture 2.1.7] (see also [1, Section 3]) on gamma-positivity, it is reasonable to expect that the generating polynomial $h(\sigma,t):=\sum_{i=0}^{n}h_{i}(\sigma)t^{i}$ of $h(\sigma)$ is $\gamma$-positive, meaning that
@@
Our second main result confirms another curious conjecture of Chapoton [5], concerning the Ehrhart polynomial of composition polytopes. We recall (see, for instance, [10, Section 4.6]) that the Ehrhart polynomial ${\rm Ehr}(Q,t)$ of a lattice polytope ${\mathcal{Q}}\subset{\mathbb{R}}_{\geq 0}^{n}$ has the defining property that ${\rm Ehr}({\mathcal{Q}},m)=\#(m{\mathcal{Q}}\cap{\mathbb{N}}^{n})$ for every $m\in{\mathbb{N}}$. The zeta polynomial (see [10, Section 3.12]) ${\mathcal{Z}}(P,t)$ of a finite poset $(P,\preceq)$ has the defining property that ${\mathcal{Z}}(P,m+1)$ is equal to the number of $m$-element multichains in $P$, meaning sequences $(p_{1},p_{2},\dots,p_{m})$ of elements of $P$ such that $p_{1}\preceq p_{2}\preceq\cdots\preceq p_{m}$, for every $m\in{\mathbb{N}}$. Given a composition $\sigma$ of $n$, Chapoton [5, Section 1] has considered the set $P_{\sigma}$ of lattice points in ${\mathcal{Q}}_{\sigma}$ as a subposet of the product poset ${\mathbb{N}}^{n}$.
@@
Theorem 1.2 is proven in Section 2; the proof yields an explicit combinatorial interpretation of the numbers $\gamma_{i}(\sigma)$ which appear in Equation (2) (see Proposition 2.2). Section 2 also shows that $h(\sigma)=h({\rm rev}(\sigma))$ for every composition $\sigma$. Two proofs of Theorem 1.3, one essentially bijective and another based on the results of [12, Section 3], are given in Section 3 and a combinatorial interpretation of the $h^{\ast}$-polynomial of ${\mathcal{Q}}_{\sigma}$ is deduced via the theory of pure shellability of posets (see Proposition 3.2). While this paper is restricted to composition polytopes, we do hope that some of the arguments presented here can be extended to more general classes of lattice polytopes.
Throughout this paper, we set ${\mathbb{N}}=\{0,1,2,\dots\}$ and $n]=\{1,2,\dots,n\}$ for $n\in{\mathbb{N}}$. We denote by #S\#S the cardinality of a finite set $S$ and write $\sigma\models n$ when $\sigma$ is a composition of $n$. Our notation and terminology about posets follows [10, Chapter 3]. Background on Ehrhart theory and face enumeration of simplicial complexes and polytopes can be found in [3, 7] [10, Chapter 4] and [9], respectively.
@@ 3. Ehrhart and Zeta polynomials
Part of our second proof of Theorem 1.3 follows the proof of [12, Equation (33)], which is a formula for the number of lattice points of a Pitman–Stanley polytope. To avoid making too long of a digression, we assume familiarity with Sections 3 and 4 of [12] and only sketch this part of the proof.
@@
Let us consider the poset $\Omega_{\sigma}$ which consists of two disjoint chains $\alpha_{1}\prec\alpha_{2}\prec\cdots\prec\alpha_{n}$ and $\beta_{1}\prec\beta_{2}\prec\cdots\prec\beta_{n}$, together with the additional cover relations $\alpha_{s_{i}}\prec\beta_{s_{i}}$ for every $i\in[k]$. We let $C=\{\beta_{1},\beta_{2},\dots,\beta_{n}\}$ and ${\mathbf{u}}=(u_{1},u_{2},\dots,u_{n})$, where $u_{j}=jm$ for every $j\in[n]$ and where $m\in{\mathbb{N}}$ is fixed. In this situation, Pitman and Stanley [12, Section 3] have defined the order cone section ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ as the polytope of all order preserving maps $f:\Omega_{\sigma}\rightarrow{\mathbb{R}}_{\geq 0}$ such that $f(\beta_{j})=u_{j}$ for every $j\in[n]$. Setting $f_{j}=f(\alpha_{j})$ for every $j\in[n]$ and $f_{0}=0$ we have
@@
for every $i\in[k]$. The transformation $x_{j}=f_{j}-f_{j-1}$ for $j\in[n]$ then shows that ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ is integrally equivalent to $m{\mathcal{Q}}_{\sigma}$. Theorem 5 in [12] expresses the number of lattice points of ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ as a sum over all linear extensions of $\Omega_{\sigma}$. Such a linear extension is uniquely determined by the positions, say $1\leq j_{1}<j_{2}<\cdots<j_{n}\leq 2n$, of $\beta_{1},\beta_{2},\dots,\beta_{n}$, which must satisfy $j_{s_{i}}\geq 2s_{i}$ for every $i\in[k]$. Thus, setting $\kappa_{i}=j_{i}-j_{i-1}-1$ for $i\in[n]$ (where $j_{0}=0$), the argument of [12, p. 617] shows that in our situation, [12, Theorem 5] yields the formula
@@
Given a word $w=(w_{1},w_{2},\dots,w_{n})\in[n]^{n}$, we denote by ${\rm asc}(w)$ the number of indices (called ascents) $i\in[n]$ such that $w_{i-1}<w_{i}$, where $w_{0}:=1$.
@@
The proof employs the basics of the theory of EL-shellability of posets [4] [13, Lecture 3], which we briefly recall. Let $(P,\preceq)$ be a finite graded poset which has a minimum element $\hat{0}$ and a maximum element $\hat{1}$ and let ${\mathcal{E}}(P)$ be the set of covering relations of $P$. An edge labeling of $P$ (with integer labels) is any map $\lambda:{\mathcal{E}}(P)\rightarrow{\mathbb{Z}}$. Let $C$ be an unrefinable chain $p_{0}\prec p_{1}\prec\cdots\prec p_{m}$ of elements of $P$, so that $(p_{i-1},p_{i})\in{\mathcal{E}}(P)$ for every $i\in[m]$. We define the label of $C$ with respect to $\lambda$ as the sequence
@@
and the descent set of $C$ as ${\rm Des}(C)=\{i\in[m-1]:\lambda(p_{i-1},p_{i})>\lambda(p_{i},p_{i+1})\}$. We say that $C$ is rising with respect to $\lambda$ if $\lambda(C)$ is weakly increasing and that $C$ is lexicographically smaller (with respect to $\lambda$) than an unrefinable chain $C^{\prime}$ in $P$ of the same length if $\lambda(C)$ precedes $\lambda(C^{\prime})$ in the lexicographic order. An edge labeling $\lambda$ of $P$ is called an EL-labeling if for every nonsingleton interval $[u,v]$ in $P$: (i) there is a unique rising (with respect to $\lambda$) maximal chain in $[u,v]$; and (ii) this chain is lexicographically smallest among all maximal chains in $[u,v]$. Then
@@
where ${\mathcal{M}}(P)$ is the set of all maximal chains of $P$ and ${\rm des}(C)=\#{\rm Des}(C)$ is the number of descents of $C$ for $C\in{\mathcal{M}}(P)$. Here $\Delta(P)$ stands for the order complex of $P$ (meaning, the simplicial complex consisting of all chains in $P$) and the $h$-polynomial of $\Delta(P)$ is defined as $h(\Delta(P),t)=\sum_{F\in\Delta(P)}t^{\#F}(1-t)^{d+1-(\#F)}$, where $d$ is the rank of $P$.
@@
Remark 3.3.
Proposition 3.2 specializes to a well known combinatorial interpretation of the $h^{\ast}$-polynomial (equivalently, of the $h$-polynomial of the $n$-fold edgewise subdivision) of the standard $n$-dimensional simplex when $\sigma=(n)$ has one part; see, for instance, [8, Corollary 8] for a generalization and refinement. A generalization of Proposition 3.2, valid for the polytopes in ${\mathbb{R}}^{n}$ defined by the equations